OpenStudy (rainstark):

Lim (1-sin3x)/(x-pi/2)^2 when x-> pi/2

OpenStudy (sooobored):

\[\sin (\frac{\pi}{2} -x)=\cos(x)\]

OpenStudy (sooobored):

or \[\cos(\frac{\pi}{2} -x) = \sin x\]

OpenStudy (rainstark):

The answer is 9/2 tho

zepdrix (zepdrix):

No, it approaches infinity. The numerator is approaching 1-(-1), and the denominator is approaching 0.